# Comparative study of magnetic exchange parameters and magnon dispersions in NiO and MnO from first principles

### v1

@@ RESULTS
Inelastic neutron scattering data for NiO and MnO 113, 114 are typically analyzed using LSWT based on the Heisenberg Hamiltonian with a fixed number of MIPs. These MIPs are fitted to the experimental data and then used to characterize the magnetic exchange interactions in these materials. The number of MIPs is chosen to ensure a good fit to the experimental magnon dispersion, while keeping the model as simple as possible. As we showed previously, from a computational standpoint, accurate convergence of the magnon dispersion in NiO and MnO requires including exchange interactions up to the fifth nearest neighbor shell (see Table 1). However, fitting experimental data with such a large number of parameters is challenging due to limited resolution and uncertainties in the measurements, and because distant interactions are small in magnitude. Therefore, experimental studies typically use a reduced model, fitting only the nearest and next-nearest neighbor interactions, as described by the simplified Heisenberg Hamiltonian 113, 114:
@@
Here, $\mathcal{J}_{1}^{-}$ and $\mathcal{J}_{1}^{+}$ are effective nearest-neighbor exchange parameters for the parallel (n.n.p) and antiparallel (n.n.a) spins, respectively, that differ due to rhombohedral lattice distortions, while $\mathcal{J}_{2}$ is the effective next-nearest-neighbor (n.n.n.) exchange parameter. These effective parameters differ from the exchange parameters $J_{1}^{-}$, $J_{1}^{+}$, and $J_{2}$ in Eq. (1) because the former ones are renormalized to account for the absence of third- to fifth-nearest-neighbor interactions in the simplified model of Eq. (2). That is, the effect of these neglected interactions are effectively absorbed into the first two shells. To clearly differentiate these fitting parameters $\mathcal{J}$ from the exchange couplings $J$ computed using MFT or $\Delta E$, we use a different font style in our notation.
@@
To fit all magnon dispersion curves in Fig. 2 using the minimal model of Eq. (2) and the respective analytical LSWT expression (see Sec. S3.3 in the SI), we employ a non-linear least-squares procedure. Importantly, we perform the fitting independently rather than relying on the fitted values reported in Refs. 113, 114, for two main reasons. First, the original works included a single-ion anisotropy term arising from magnetocrystalline effects, which we neglect in Eq. (2). While this term only slightly renormalizes $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$, we omit it to ensure a consistent comparison focused solely on the first- and second-nearest neighbor exchange interactions, avoiding additional sources of discrepancy between theory and experiment. Second, we adopt spin quantum numbers $S=1$ for Ni2+ and $S=5/2$ for Mn2+, while Ref. 114 used $S=2.445$ for Mn2+. Although this is a minor difference, we use standard spin values to ensure maximal consistency with our computational models and to enable an accurate comparison of the fitted exchange parameters. Finally, for NiO, the magnon energy at the M point in Fig. 2 is essential for a reliable fit due to the limited number of experimental points along the $\Gamma$–M path in the BZ. We use the value 4.54 meV, consistent with Ref. 113, which was based on infrared measurements from Refs. 115, 116.
@@
The fitting curves to the experimental and TDDFPT+$U$ magnon dispersions are shown in Fig. 2, while the corresponding fits of the LSWT-derived magnon dispersions using the $J_{1}$$J_{5}$ exchange parameters from MFT and $\Delta E$ (listed in Table 1) are presented in Sec. S2 of the SI. As seen in Fig. 2, the TDDFPT+$U$ data are fitted with high accuracy. The fit to the experimental data is also satisfactory, considering the substantial spread in the measurements caused by the limited resolution of inelastic neutron scattering. Similarly, the fitted curves show excellent agreement with the LSWT dispersions based on the MFT and $\Delta E$ parameters (see Sec. S2 of the SI). This confirms that the minimal Heisenberg model in Eq. (2) is sufficient to reproduce all the features of the magnon dispersions in NiO and MnO with high accuracy. In the following, we compare the fitted exchange parameters $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ with those computed from first principles using the MFT and $\Delta E$ methods, namely $J_{1}^{-}$, $J_{1}^{+}$, and $J_{2}$.
@@
Figure 3 summarizes the effective exchange parameters $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ for NiO, obtained by fitting the magnon dispersions from TDDFPT+$U$ and LSWT-derived ones using the $J_{1}$$J_{5}$ values from MFT and $\Delta E$, and compare them to those fitted to experimental data 113, 114. The MFT-fitted parameters deviate most from experiment, while the $\Delta E$-fitted parameters are in closest agreement. TDDFPT+$U$-fitted parameters yield intermediate accuracy, slightly less accurate than the $\Delta E$-fitted values. Despite differences in accuracy, all methods consistently reproduce the trend $|\mathcal{J}_{1}^{+}|<|\mathcal{J}_{1}^{-}|\ll|\mathcal{J}_{2}|$. Notably, the trend $|\mathcal{J}_{1}^{+}|<|\mathcal{J}_{1}^{-}|$ is crucial for ensuring positive-definiteness of the magnon spectrum in the minimal Heisenberg model (2). We recall that this trend is reversed in the raw $\Delta E$ data for NiO [see Table 1], leading to a magnon instability near the M point [see Fig. S1(a) in the SI]. The comparison of the fitted effective ($\mathcal{J}$) and computed ($J$) parameters reveals systematic renormalizations: The average deviation for MFT and $\Delta E$ is $\Delta(\mathcal{J}_{1}^{+}-J_{1}^{+})$ is 19% and $\Delta(\mathcal{J}_{1}^{-}-J_{1}^{-})$ is 7%, while $\Delta(\mathcal{J}_{2}-J_{2})$ is 2% for MFT and 10% for $\Delta E$. These differences underscore how excluding third- to fifth-nearest-neighbor terms and absorbing their effects into renormalized $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ in the minimal model influences the resulting magnon dispersion in NiO.
@@
Figure 4 shows the effective fitted exchange parameters $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ for MnO. As in NiO, MFT-fitted values deviate most from experiment, $\Delta E$-fitted values show the best agreement, and TDDFPT+$U$ yields intermediate accuracy. However, unlike NiO, the fitted parameter trends differ across methods: MFT gives $|\mathcal{J}_{1}^{+}|>|\mathcal{J}_{2}|>|\mathcal{J}_{1}^{-}|$, $\Delta E$ yields $|\mathcal{J}_{1}^{+}|=|\mathcal{J}_{2}|>|\mathcal{J}_{1}^{-}|$, and TDDFPT+$U$ gives $|\mathcal{J}_{2}|>|\mathcal{J}_{1}^{+}|>|\mathcal{J}_{1}^{-}|$. Only the TDDFPT+$U$ trend matches the experimental trend. Moreover, the MFT-fitted trend matches the trend for computed $J$ values using MFT, while the $\Delta E$-fitted trend differs from its corresponding trend for computed $J$ values using $\Delta E$. Unlike in NiO, these variations in trends for MnO do not affect magnon stability. The comparison of effective ($\mathcal{J}$) and computed ($J$) parameters reveals renormalization effects in MnO that differ in magnitude from those observed in NiO. More specifically, for MnO, $\Delta(\mathcal{J}_{1}^{+}-J_{1}^{+})$ is 3% for MFT and 0% for $\Delta E$, which are much smaller than in NiO; $\Delta(\mathcal{J}_{1}^{-}-J_{1}^{-})$ is 7% for both methods, similarly to NiO; $\Delta(\mathcal{J}_{2}-J_{2})$ is 6% for MFT and 20% for $\Delta E$, significantly larger than in NiO. These differences highlight the strong material dependence of the exchange parameter renormalization in the minimal Heisenberg model.
DISCUSSION
@@
Accurate modeling of magnetic excitations in transition-metal compounds remains a challenge for DFT-based approaches. In this work, we employed a consistent DFT+$U$ ground state for all calculations, with the Hubbard $U$ parameter computed from first principles via DFPT 86. This eliminates ambiguities related to empirical $U$ choices and allows a focused evaluation of each method’s accuracy in computing MIPs and magnon dispersions based on their underlying physical approximations.
@@
First, we showed that direct methods for extracting exchange parameters require careful convergence tests with respect to the interaction range in the Heisenberg Hamiltonian (1). For both NiO and MnO, we found that including exchange interactions up to the fifth nearest neighbor is needed when using the MFT or $\Delta E$ methods; otherwise, unphysical features such as magnon instabilities may arise (see e.g. the $\Delta E$-derived magnon dispersion of NiO in Fig. S1 in the SI). Additionally, we demonstrated the importance of including the rhombohedral lattice distortion, which splits the nearest-neighbor interaction into two distinct parameters, $J_{1}^{+}$ and $J_{1}^{-}$, due to magnetoelastic coupling. This splitting is crucial for reproducing the finite magnon energy at the M point in the BZ.
@@
These findings naturally raise important questions: Why do LSWT with $\Delta E$-derived exchange parameters and TDDFPT+$U$ give the best match with experiment? And why does MFT perform worst? In the following, we briefly analyze the fundamental differences between these approaches to shed light on these points.
The key features of TDDFPT+$U$ can be summarized as follows: $(i)$ Direct calculation of magnon dispersions: TDDFPT+$U$ computes magnon spectra without relying on a low-energy model, such as the Heisenberg Hamiltonian in combination with LSWT. It includes magnetic pairwise exchange interactions to infinite order; i.e., inherently accounting for all neighbor interactions (e.g., $J_{3}$, $J_{4}$, $J_{5}$, and beyond) without any truncation. $(ii)$ Fully dynamical formalism: Unlike static methods such as MFT and $\Delta E$, TDDFPT+$U$ operates in the frequency domain by solving time-dependent linear-response equations. This allows direct access to the dynamical spin susceptibility, including Hubbard $U$ corrections 39, thereby extending the DFT+$U$ framework to excited states. In contrast, dynamics in MFT and $\Delta E$ enters only through LSWT applied to a static spin Hamiltonian. $(iii)$ Coupled spin-charge response: TDDFPT+$U$ treats the linear response of both charge and magnetization densities on equal footing, naturally including their coupling. This avoids the adiabatic decoupling assumption inherent in MFT and $\Delta E$, resulting in a more consistent description of magnetic excitations. Given these advantages, one might expect TDDFPT+$U$ to yield the most accurate magnon dispersions. Yet, we find that LSWT with $\Delta E$-derived exchange parameters provides slightly better agreement with experimental data. This observation highlights the need for further benchmarking these methods across a broader range of materials to better understand their relative strengths and limitations, but also serves as a reminder of the limitations of even the state-of-the-art in energy functionals (hence, the slightly better agreement might be accidental).
We now turn to a more detailed discussion of the $\Delta E$ and MFT methods. Starting with the $\Delta E$ approach: since it computes exchange parameters from total-energy differences between magnetic configurations, accurately resolving very small differences, such as the subtle splitting between nearest-neighbor couplings caused by weak rhombohedral distortion (e.g., $|\mathcal{J}_{1}^{+}-\mathcal{J}_{1}^{-}|\sim 0.01$ meV in NiO), requires stringent numerical convergence of DFT+$U$ calculations for large supercells. Although the total energies involved are several orders of magnitude larger than these differences, modern numerical algorithms can achieve the necessary precision. This is analogous, e.g., to the reliable calculation of phonon frequencies using frozen-phonon methods based on supercells and finite differences 117. However, achieving high precision convergence in large supercells can be challenging for some systems and computationally demanding, especially as system sizes grow. Despite these challenges, we find that the $\Delta E$ method performs remarkably well, accurately reproducing both exchange parameters and magnon dispersions, provided that the Heisenberg model includes a sufficient number of neighboring shells to ensure convergence.
In contrast, the MFT method produces the least accurate results in our study. One of the key difference between the $\Delta E$ and TDDFPT+$U$ methods on one hand, and MFT on the other, lies in their treatment of the screening of charge and magnetization densities. Both the $\Delta E$ and TDDFPT+$U$ approaches explicitly incorporate electronic screening effects, whereas MFT neglects them. Specifically, the $\Delta E$ method, as a finite-difference scheme, accounts for screening via self-consistent electronic ground-state minimizations of different spin configurations 26, 27, 28, 29. The TDDFPT+$U$ method evaluates the dynamical response of charge and magnetization densities to linear order in the external magnetic perturbation 71, 39. In contrast, MFT is a non-self-consistent approach that approximates the screened perturbing potential by the bare external potential, thereby ignoring electronic screening effects 50, 51. If MFT were extended to include variations of the Kohn-Sham potential induced by infinitesimal spin rotations, its accuracy might approach that of the $\Delta E$ and TDDFPT+$U$ methods.
In addition, the lowest accuracy of MFT could be attributed in part to the specific implementation employed, via the TB2J code 51, which introduces some approximations. As previously discussed, MIPs obtained via MFT are highly sensitive to technical details of the implementation 50, 49, 54, 55. In particular, Ref. 49 points out concerns stemming from specific algorithmic choices (we refer the reader to Sec.VII of Ref.49 for a detailed discussion). A full analysis of these issues lies beyond the scope of our work, as it would require a deeper examination of the theoretical foundations of MFT. Moreover, recent studies 54, 55 highlight that the choice of variables used to represent spin rotations can substantially affect the resulting MIPs, which is another source of variability in MFT-based calculations. These findings underscore the need for further investigation using alternative MFT implementations to better evaluate the method’s accuracy and reliability.

### v2

@@ III.4 Fitting effective exchange parameters of the minimal Heisenberg model
Inelastic neutron scattering data for NiO and MnO [130, 131] are typically analyzed using LSWT based on the Heisenberg Hamiltonian with a fixed number of MIPs. These MIPs are fitted to the experimental data and then used to characterize the magnetic exchange interactions in these materials. The number of MIPs is chosen to ensure a good fit to the experimental magnon dispersion, while keeping the model as simple as possible. As we showed previously, from a computational standpoint, accurate convergence of the magnon dispersion in NiO and MnO requires including exchange interactions up to the fifth nearest neighbor shell (see Table 1). However, fitting experimental data with such a large number of parameters is challenging due to limited resolution and uncertainties in the measurements, and because distant interactions are small in magnitude. Therefore, experimental studies typically use a reduced model, fitting only the nearest and next-nearest neighbor interactions, as described by the simplified Heisenberg Hamiltonian [130, 131]:
@@
Here, $\mathcal{J}_{1}^{-}$ and $\mathcal{J}_{1}^{+}$ are effective nearest-neighbor exchange parameters for the parallel (n.n.p) and antiparallel (n.n.a) spins, respectively, that differ due to rhombohedral lattice distortions, while $\mathcal{J}_{2}$ is the effective next-nearest-neighbor (n.n.n.) exchange parameter. These effective parameters differ from the exchange parameters $J_{1}^{-}$, $J_{1}^{+}$, and $J_{2}$ in Eq. (1) because the former ones are renormalized to account for the absence of third- to fifth-nearest-neighbor interactions in the simplified model of Eq. (2). That is, the effect of these neglected interactions is effectively absorbed into the first two shells. To clearly differentiate these fitting parameters $\mathcal{J}$ from the exchange couplings $J$ computed using MFT or $\Delta E$, we use a different font style in our notation.
@@
To fit all magnon dispersion curves in Fig. 2 using the minimal model of Eq. (2) and the respective analytical LSWT expression (see Sec. S3 in the SM [115]), we employ a non-linear least-squares procedure. Importantly, we perform the fitting independently rather than relying on the fitted values reported in Refs. 130, 131, for two main reasons. First, the original works included a single-ion anisotropy term arising from magnetocrystalline effects, which we neglect in Eq. (2). We verified that this term only slightly renormalizes $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ (at most by a few percent), and therefore we omit it to ensure a consistent comparison focused solely on the first- and second-nearest neighbor exchange interactions. Second, we adopt spin quantum numbers $S=1$ for Ni2+ and $S=5/2$ for Mn2+, while Ref. 131 used $S=2.445$ for Mn2+. Although this is a minor difference, we use nominal spin values to ensure maximal consistency with our computational models and to enable an accurate comparison of the fitted exchange parameters. Finally, for NiO, the magnon energy at the M point in Fig. 2 is essential for a reliable fit due to the limited number of experimental points along the $\Gamma$–M path in the BZ. We use the value 4.54 meV, consistent with Ref. 130, which was based on infrared measurements from Refs. 132, 133.
@@
The fitting curves to the experimental and TDDFPT+$U$ magnon dispersions are shown in Fig. 2, while the corresponding fits of the LSWT-derived magnon dispersions using the $J_{1}$$J_{5}$ exchange parameters from MFT and $\Delta E$ (listed in Table 1) are presented in Sec. S2 of the SM [115]. As seen in Fig. 2, the TDDFPT+$U$ data are fitted with high accuracy. The fit to the experimental data is also satisfactory, considering the substantial spread in the measurements caused by the limited resolution of inelastic neutron scattering. Similarly, the fitted curves show excellent agreement with the LSWT dispersions based on the MFT and $\Delta E$ parameters (see Sec. S2 of the SM [115]). This confirms that the minimal Heisenberg model in Eq. (2) is sufficient to reproduce all the features of the magnon dispersions in NiO and MnO with high accuracy. In the following, we compare the fitted exchange parameters $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ with those computed from first principles using the MFT and $\Delta E$ methods, namely $J_{1}^{-}$, $J_{1}^{+}$, and $J_{2}$.
@@
Figure 3 summarizes the effective exchange parameters $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ for NiO, obtained by fitting the magnon dispersions from TDDFPT+$U$ and LSWT-derived ones using the $J_{1}$$J_{5}$ values from MFT and $\Delta E$, and compare them to those fitted to experimental data [130, 131]. The MFT-fitted parameters deviate most from experiment, while the $\Delta E$-fitted parameters are in closest agreement. TDDFPT+$U$-fitted parameters yield intermediate accuracy, slightly less accurate than the $\Delta E$-fitted values. Despite differences in accuracy, all methods consistently reproduce the trend $|\mathcal{J}_{1}^{+}|<|\mathcal{J}_{1}^{-}|\ll|\mathcal{J}_{2}|$. Notably, the trend $|\mathcal{J}_{1}^{+}|<|\mathcal{J}_{1}^{-}|$ is crucial for ensuring positive-definiteness of the magnon spectrum in the minimal Heisenberg model (2). We recall that this trend is reversed in the raw $\Delta E$ data for NiO (see Table 1), leading to a magnon instability near the M point [see Fig. S1(a) in the SM [115]]. The comparison of the fitted effective ($\mathcal{J}$) and computed ($J$) parameters reveals systematic renormalizations: The average deviation $\Delta(\mathcal{J}_{1}^{+}-J_{1}^{+})$ is 14% and $\Delta(\mathcal{J}_{1}^{-}-J_{1}^{-})$ is 10% for both MFT and $\Delta E$, while $\Delta(\mathcal{J}_{2}-J_{2})$ is 2% for MFT and 14% for $\Delta E$. These differences underscore how excluding third- to fifth-nearest-neighbor terms and absorbing their effects into renormalized $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ in the minimal model influences the resulting magnon dispersion in NiO.
@@
Figure 4 shows the effective fitted exchange parameters $\mathcal{J}_{1}^{-}$, $\mathcal{J}_{1}^{+}$, and $\mathcal{J}_{2}$ for MnO. As in NiO, MFT-fitted values deviate most from experiment, $\Delta E$-fitted values show the best agreement, and TDDFPT+$U$ yields intermediate accuracy. However, unlike NiO, the fitted parameter trends differ across methods: MFT and $\Delta E$ give $|\mathcal{J}_{1}^{+}|>|\mathcal{J}_{2}|>|\mathcal{J}_{1}^{-}|$, while TDDFPT+$U$ gives $|\mathcal{J}_{2}|>|\mathcal{J}_{1}^{+}|>|\mathcal{J}_{1}^{-}|$. Only the TDDFPT+$U$ trend matches the experimental trend. Moreover, the MFT-fitted trend matches the trend for computed $J$ values using MFT, while the $\Delta E$-fitted trend differs from its corresponding trend for computed $J$ values using $\Delta E$. Unlike in NiO, these variations in trends for MnO do not affect magnon stability. The comparison of effective ($\mathcal{J}$) and computed ($J$) parameters reveals renormalization effects in MnO that differ in magnitude from those observed in NiO. More specifically, for MnO, the average $\Delta(\mathcal{J}_{1}^{+}-J_{1}^{+})$ is 5% for MFT and $\Delta E$, which is 3 times smaller than in NiO; the average $\Delta(\mathcal{J}_{1}^{-}-J_{1}^{-})$ is also 5% for both methods, which is 2 times smaller than in NiO; $\Delta(\mathcal{J}_{2}-J_{2})$ is 6% for MFT and 22% for $\Delta E$, which is 232-3 times larger than in NiO. These differences highlight the strong material dependence of the exchange parameter renormalization in the minimal Heisenberg model.
It is useful to comment on the trend observed in the renormalization of exchange parameters when the magnon dispersions derived from the full set of ab initio couplings are fitted using the minimal Heisenberg model. For both NiO and MnO, the effective fitted parameters are generally smaller than the computed ones, which reflects how the longer-range interactions (when omitted explicitly) are effectively absorbed into the nearest- and next-nearest-neighbour terms. Their signs and relative magnitudes influence the curvature and overall dispersion in such a way that the minimal model reproduces the correct magnon energies only by adjusting (renormalizing) the retained couplings. Whether similar trends hold more broadly remains an open question and will require further systematic studies across different classes of magnetic materials.
IV DISCUSSION
The findings of this study naturally raise important questions: Why do LSWT with $\Delta E$-derived exchange parameters and TDDFPT+$U$ give the best match with experiment? And why does MFT perform worst? In the following, we briefly analyze the fundamental differences between these approaches to shed light on these points.
The key features of TDDFPT+$U$ can be summarized as follows: $(i)$ Direct calculation of magnon dispersions: TDDFPT+$U$ computes magnon spectra without relying on a low-energy model, such as the Heisenberg Hamiltonian in combination with LSWT. It includes magnetic pairwise exchange interactions to infinite order; i.e., inherently accounting for all neighbor interactions (e.g., $J_{3}$, $J_{4}$, $J_{5}$, and beyond) without any truncation. $(ii)$ Fully dynamical formalism: Unlike static methods such as MFT and $\Delta E$, TDDFPT+$U$ operates in the frequency domain by solving time-dependent linear-response equations. This allows direct access to the dynamical spin susceptibility, including Hubbard $U$ corrections [39], thereby extending the DFT+$U$ framework to excited states. In contrast, dynamics in MFT and $\Delta E$ enters only through LSWT applied to a static spin Hamiltonian. $(iii)$ Coupled spin-charge response: TDDFPT+$U$ treats the linear response of both charge and magnetization densities on equal footing, naturally including their coupling. This avoids the adiabatic decoupling assumption inherent in MFT and $\Delta E$, resulting in a more consistent description of magnetic excitations. Given these advantages, one might expect TDDFPT+$U$ to yield the most accurate magnon dispersions. Yet, we find that LSWT with $\Delta E$-derived exchange parameters provides slightly better agreement with experimental data. This observation highlights the need for further benchmarking these methods across a broader range of materials to better understand their relative strengths and limitations, but also serves as a reminder of the limitations of the employed energy functionals (hence, the slightly better agreement might be accidental). In particular, using more advanced functionals, such as the extended Hubbard functionals (DFT+UU+VV[134], which are especially relevant for systems with significant covalent interactions [90, 91, 92, 93, 94, 135, 95], or the orbital-resolved DFT+$U$ formulation [99, 136], could change this trend. Specifically, TDDFPT+UU+$V$ or orbital-resolved TDDFPT+$U$ might ultimately outperform LSWT with $\Delta E$-derived parameters based on DFT+UU+$V$ or orbital-resolved DFT+$U$, respectively.
We now turn to a more detailed discussion of the $\Delta E$ and MFT methods. Starting with the $\Delta E$ approach: since it computes exchange parameters from total-energy differences between magnetic configurations, accurately resolving very small differences, such as the subtle splitting between nearest-neighbor couplings caused by weak rhombohedral distortion (e.g., $|\mathcal{J}_{1}^{+}-\mathcal{J}_{1}^{-}|\sim 0.01$ meV in NiO), requires stringent numerical convergence of DFT+$U$ calculations for large supercells. Although the total energies involved are several orders of magnitude larger than these differences, modern numerical algorithms can achieve the necessary precision. This is analogous, e.g., to the reliable calculation of phonon frequencies using frozen-phonon methods based on supercells and finite differences [137]. However, achieving high precision convergence in large supercells can be challenging for some systems and computationally demanding, especially as system sizes grow. Regarding this point, our results for NiO and MnO clearly indicate that including exchange interactions up to the fifth nearest neighbor is essential for converged and stable magnon dispersions. However, performing $\Delta E$ calculations for such long-range couplings requires very large supercells containing several hundred atoms, and the associated computational cost scales cubically with system size. This challenge becomes even more significant in materials where long-range and anisotropically split interactions play a key role, such as altermagnets [12, 13]. In these systems, the anisotropic splitting of distant exchange shells must be resolved, for instance, the 7th-nearest-neighbor exchange splitting in MnF2 [138] or even the 10th-nearest-neighbor interactions in MnTe [139]. Converging such distant exchange interactions requires very large supercells and high numerical cost, which can make the $\Delta E$ approach increasingly difficult to apply. Despite these challenges, we find that the $\Delta E$ method performs remarkably well in NiO and MnO, accurately reproducing both exchange parameters and magnon dispersions, provided that the Heisenberg model includes a sufficient number of neighboring shells to ensure convergence.
In contrast, the MFT method produces the least accurate results in our study. One of the key differences between the $\Delta E$ and TDDFPT+$U$ methods on one hand, and MFT on the other, lies in their treatment of the screening of charge and magnetization densities. Both the $\Delta E$ and TDDFPT+$U$ approaches explicitly incorporate electronic screening effects, whereas MFT neglects them. Specifically, the $\Delta E$ method, as a finite-difference scheme, accounts for screening via self-consistent electronic ground-state minimizations of different spin configurations [26, 27, 28, 29]. The TDDFPT+$U$ method evaluates the dynamical response of charge and magnetization densities fully self-consistently to linear order in the external magnetic perturbation [71, 39]. In contrast, MFT is a non-self-consistent approach that approximates the screened perturbing potential by the bare external potential, thereby ignoring electronic screening effects [50, 51]. In this context, we note the recent work of Ref. [140], which compared self-consistent and non-self-consistent formulations of the spin-spiral method. It was shown that neglecting self-consistency leads to substantial deviations in MIPs and magnon energies relative to TDDFT benchmark data for elemental $3d$ metallic ferromagnets (Fe, Co, and Ni) and selected Heusler compounds. These findings are in line with our observations that the MFT results diverge markedly from those obtained using the self-consistent $\Delta E$ and TDDFPT+$U$ approaches. Therefore, if MFT were extended to include variations of the Kohn-Sham potential induced by infinitesimal spin rotations, its accuracy might approach that of the $\Delta E$ and TDDFPT+$U$ methods. However, it is important to note that the observed deviations of MFT are material-dependent, particularly on the magnitude of the magnetic moments of TM ions and the degree of electron localization versus itinerancy [140]. Hence, a more systematic investigation across a larger and more diverse set of materials is needed to fully assess the accuracy of the MFT method.
In addition, the lowest accuracy of MFT could be attributed in part to the specific implementation employed, via the TB2J code [51], which introduces some approximations. As previously discussed, MIPs obtained via MFT are highly sensitive to technical details of the implementation [50, 49, 54, 55]. In particular, Ref. 49 points out concerns stemming from specific algorithmic choices (we refer the reader to Sec.VII of Ref. 49 for a detailed discussion). A full analysis of these issues lies beyond the scope of our work, as it would require a deeper examination of the theoretical foundations of MFT. Moreover, recent studies [54, 55] highlight that the choice of variables used to represent spin rotations can substantially affect the resulting MIPs, which is another source of variability in MFT-based calculations. Furthermore, as noted earlier, the accuracy of MFT implemented in TB2J for plane-wave-based codes is extremely sensitive to the quality of the wannierization, particularly the atom-centering of the resulting Wannier functions. Hence, to further assess the accuracy of MFT, alternative wannierization strategies could be explored, such as symmetry-adapted Wannier functions [122] or even approaches that avoid Wannier functions altogether, as in the recently proposed method of Ref. [141]. Finally, using alternative MFT implementations, for instance based on KKR [42, 43] or LMTO [44, 45], could provide further insights into the accuracy and reliability of the MFT method. However, a major challenge is that these approaches rely on different DFT codes and use distinct levels of theory for the ground-state description, which differ from ours and complicate the direct comparison of computed MIPs and magnon dispersions [24, 105].
V Conclusions
@@
Accurate modeling of magnetic excitations in transition-metal compounds remains a challenge for DFT-based approaches. In this work, we employed a consistent DFT+$U$ ground state for all calculations, with the Hubbard $U$ parameter computed from first principles via DFPT [86]. This eliminates ambiguities related to empirical $U$ choices and allows a focused evaluation of each method’s accuracy in computing MIPs and magnon dispersions based on their underlying physical approximations.
@@
First, we showed that direct methods for extracting exchange parameters require careful convergence tests with respect to the interaction range in the Heisenberg Hamiltonian (1). For both NiO and MnO, we found that including exchange interactions up to the fifth nearest neighbor is needed when using the MFT or $\Delta E$ methods; otherwise, unphysical features such as magnon instabilities may arise (see e.g. the $\Delta E$-derived magnon dispersion of NiO in Fig. S1 in the SM [115]). Additionally, we demonstrated the importance of including the rhombohedral lattice distortion, which splits the nearest-neighbor interaction into two distinct parameters, $J_{1}^{+}$ and $J_{1}^{-}$, due to magnetoelastic coupling. This splitting is crucial for reproducing the finite magnon energy at the M point in the BZ.